# Algorithmic Collusion And The Minimum Price Markov Game

### v1

@@ 4.1 The Minimum Price Game: A Single-Stage Formulation
Strategies. All players are symmetric in their strategy set: for each firm $i$, $s_{i}=\{FP,CP\}$, where Fair Price (FP) is associated with bidding the fair price $b_{i}$ stemming from Equation (1), and Collusive Price (CP) with bidding a higher price. The collusive bid is defined by $\alpha\cdot b_{i}$, where $\tau\geq\alpha>1$, with the upper bound $\tau$ ensuring realistic bids. We denote the strategy profile for a given occurrence as the tuple $(s_{1},s_{2},\dots,s_{n})$. In the static setting, the specific bid values as well as the actual contract value are illustrative and not numerically significant since the payoffs are directly tied to the strategy profiles played.
Payoffs. Let $u_{i}(FP,k)$ be the payoff of firm $i$ if they bid the fair price while $k$ competitors bid the fair price, and $u_{i}(CP,k)$ the payoff of firm $i$ if they bid the collusive price while $k$ competitors bid the fair price. The total market power of lowest-bidding firms is
$\beta_{\Omega}=\sum_{j\in\Omega}\beta_{j}.$ (2)
The minimum price rule fosters a tension between group and individual rationality, and this dilemma must be represented in the payoff structure. The key idea is that coordination must be unanimous to achieve a collusive spoil, meaning that $u_{i}(CP,k=0)>u_{i}(FP,k=n-1)$. However, when $k=0$, that is, when all competitors bid the collusive price, $u_{i}(PF,k)>u_{i}(CP,k)$. Having in mind the fact that payoffs are shared in case of ties (following Assumption 2), and using Equation (2), we define the resulting payoff for player $i$ when they bid the fair price while $k=0$ as
$u_{i}(FP,k)=\frac{\beta_{i}}{\beta_{\Omega}}\cdot\alpha\cdot b_{i}.$ (3)
Conversely, when $k>0$, that is, if at least one competitor bid the fair price, $u_{i}(PF,k)>u_{i}(CP,k)$. Again following Assumption 2, and using Equation (2), we define the payoff for the player $i$ bidding the fair price while $k>0$ as
$u_{i}(FP,k)=\frac{\beta_{i}}{\beta_{\Omega}}\cdot b_{i}.$ (4)
In the static setting, actual payoff values $u_{i}(PF,k)$ when $k=0$ and $u_{i}(CP,k)$ when $k>0$ are not crucial as long as the payoff inequalities are respected.
Analysis. Equations (2), (3), and (4) define the payoff structure in the general heterogeneous case where the market power $\beta_{i}$ differs among firms. A perfectly homogeneous market is a special case where all players share the same market power, meaning $\beta_{i}=\frac{1}{n}$ for all $i\in\mathcal{N}$. In such scenarios, when collusion is achieved (i.e., player $i$ bids the collusive price and $k=0$), Equation (2) becomes $\beta_{\Omega}=1$ since $|\Omega|=n$, and hence Equation (3) becomes $u_{i}(CP,k)=\frac{1}{n}\cdot\alpha\cdot b_{i}$. Similarly, when player $i$ bids the fair price alongside at least one competitor (i.e., $k>0$), Equation (4) becomes $u_{i}(FP,k)=\frac{1}{(k+1)}\cdot b_{i}$ for all $i\in\Omega$.
@@
The 2-player homogeneous Minimum Price Game (MPG), i.e., when both players share the same market power, as derived from Assumptions 1 to 4, is a 2-player classical Prisoner’s Dilemma.
@@
Suppose $n=2$, the game being composed of player $i$ and its opponent player $o$, both sharing the market equally ($\beta_{i}=\beta_{o}=0.5$). According to the payoff structure of the MPG, we can define each strategy profile for the 2-player case as follows:
  • $T$: Temptation payoff (bids the fair price while the other does not),
  • $R$: Reward payoff (both bid the collusive price),
  • $P$: Punishment payoff (both bid the fair price),
  • $S$: Sucker’s payoff (bids the collusive price while the other does not).
The 2-player homogeneous MPG is a classic Prisoner’s Dilemma as it respects the following three sufficient conditions:
  1. 1.
    $(FP,FP)$ is the dominant strategy profile.
    The minimum price rule inherently favors the lowest bidder. Thus, if player $o$ bids the fair price, the payoffs are
    $u_{i}(FP,FP)=\frac{1}{2}\cdot b_{i}=P\quad\text{and}\quad u_{i}(CP,FP)=0=S,\quad\text{therefore}\quad P>S.$
    In addition, since the players are symmetrical, $u_{i}=u_{o}$ for any strategy profile $(s_{i},s_{o})$. This means that bidding the fair price is the best response for both players, regardless of their opponent’s strategy. Thus, the strategy profile $(FP,FP)$ is the unique Nash equilibrium.
  2. 2.
    The strategy profile $(CP,CP)$ is Pareto dominant.
    A strategy profile is Pareto-efficient if no other strategy can make one player better off without making the other worse off. Since
    $u_{i}(CP,CP)=\frac{1}{2}\cdot\alpha\cdot b_{i}>u_{i}(FP,FP)=\frac{1}{2}\cdot b_{i}\quad\text{because}\quad\alpha>1,$
    $R>P$, meaning that $(CP,CP)$ is Pareto-efficient.
    A strategy profile is Pareto dominant if it is Pareto-efficient, and there is no other Pareto-efficient strategy profile that provides a higher payoff for at least one player without reducing the payoff for the other. The profiles $(CP,FP)$ and $(FP,CP)$ are also Pareto-efficient because the resulting bi-matrix game is symmetric, i.e,
    $u_{i}(CP,FP)=u_{o}(FP,CP)=S.$
    Since $R>P$ and $R>S$, the strategy profile $(CP,CP)$ is Pareto dominant.
  3. 3.
    The payoff structure respects the inequality $T>R>P>S$.
    We have already shown that $R>P>S$. Since
    $u_{i}(FP,CP)=b_{i}>u_{i}(CP,CP),$
    $T>R$, and therefore, $T>R>P>S$.
@@
The n-player homogeneous MPG is a n-player prisoner’s dilemma.
@@
The 2-player MPG is a 2-player Prisoner’s Dilemma (Proposition 1). It is well established in game theory that the characteristics of the 2-player Prisoner’s Dilemma generalize to the $n$-player case. Therefore, the $n$-player homogeneous MPG also exhibits the same properties, making it an $n$-player Prisoner’s Dilemma. ∎
Therefore, the homogeneous MPG provides a simple yet profound framework for the study of strategic interactions among bidders in the context of minimum price-ruled auction markets. Mutual defection leads to suboptimal outcomes for all parties involved, but aligns with the prediction of market equilibrium. If firms were to cooperatively opt for higher prices, the group outcomes would be improved, highlighting the tension between competitive strategies and potential cooperative gains.
In the heterogeneous case, however, the market power $\beta$ differs among firms. This disparity leads to a variation in the payoffs. For a low-market power firm, bidding the fair price becomes more attractive because the payoff $u_{i}(PF,CP)$ can exceed the payoff $u_{i}(CP,CP)$ when all cooperate. The proportion of the reward $\frac{\beta_{i}}{\beta_{\Omega}}$ determines each firm’s payoff, potentially resulting in higher rewards for firms with larger market shares. Consequently, the strategy profile $(CP,CP)$ is no longer Pareto dominant because the incentive to bid the fair price is stronger for firms with lower market power, and the heterogeneous MPG is hence not a classical Prisoner’s Dilemma.
@@ 4.2 Markov Game Formulation
Action space. The action space $\mathcal{A}=\{0,1\}$ is binary, reflecting the two strategies available to players, with bidding the fair price being represented by 00 and bidding the collusive price with $1$. This representation simplifies the computation of action and joint action frequencies. Let $a_{i}^{t}\in\mathcal{A}$ be the action taken by agent $i$ at time $t$, and $f^{t}\in\mathcal{A}^{n}$ be the joint action $(a_{1},\dots,a_{n})$ in auction $t$. The action and joint action frequencies up to iteration $t$ are then respectively given by
@@
State space. The choice of state space in a Markov game is often at the discretion of the implementer, depending on the specifics of the problem and the goals of the implementation. However, for our purposes, we define the MPMG with core variables that capture essential aspects of the game. Specifically, these sets include all players’ information on action frequencies $\mathbf{\bar{a}}^{t}=\{\bar{a}_{j}^{t}\}_{j\in\mathcal{N}}$, joint action frequencies $\mathbf{\bar{f}}^{t}=\{\bar{f}_{j}^{t}\}_{j\in\mathcal{N}}$, and power parameters $\mathbf{\beta}=\{\beta_{j}\}_{j\in\mathcal{N}}$. Therefore, the state space at any time $t$ is $\mathcal{S}=\{\mathbf{\bar{a}}^{t},\mathbf{\bar{f}}^{t},\mathbf{\beta}\}$, and is of size $2n+|\mathcal{A}|^{n}$. This minimal configuration for the state space aligns with Assumption 5. Furthermore, the power parameters also allow for the possibility of modeling a dynamic market in which firms could lose or gain market power.
Reward function. The payoff structure defined for the MPG translates to a reward function in the MPMG, emphasizing the inherent connection between the Markov game and its RL implementation. The reward function, together with the transition function, which defines how the game state evolves in response to actions, fundamentally shapes the dynamics and outcomes of the game. In fact, in the classical RL setting, the normalized reward function for firm $i$ at each iteration $t$ is expressed as:
@@
where $\gamma$ is the discount factor and $r_{i}^{t-m}$ is the reward received by firm $i$ at time $t-m$. The MPMG relies on the tension between Nash and Pareto dominant strategies inherent in the classical Prisoner’s Dilemma. However, as with the state space, the reward function can be adjusted for specific implementation purposes (e.g., reward shaping) as long as the fundamental dynamics are preserved. Following Assumption 4, we can further assume that the contract value $v$ remains equivalent as the iterations unfold, normalizing them to $1$. Equation (1) then becomes
@@
We do not make this assumption formal, as this normalization does not affect the generality of the model but only simplifies the calculations. The resulting bid $b_{i}$ can also be normalized to express a fair bid of $1$ and a collusive bid of $\alpha\cdot b_{i}=\alpha$, where $\tau\geq\alpha>1$. Considering these numerical adjustments, we get the immediate reward for player $i$ at any auction $t$ in the homogeneous case as
$r_{i}=\begin{cases}\frac{1}{(k+1)}&\text{if }a_{i}=0\\ \frac{\alpha}{n}&\text{if }a_{i}=1\text{ and }k=0\\ 0&\text{if }a_{i}=1\text{ and }k>0\end{cases}.$
Note that for simplicity, we dropped the subscript $t$. When firms are heterogeneous, the immediate rewards are weighted by their market power, as in Equation (3) and (4):
$r_{i}=\begin{cases}\frac{\beta_{i}}{\beta_{\Omega}}\cdot\frac{1}{(k+1)}&\text{if }a_{i}=0\\ \frac{\beta_{i}}{\beta_{\Omega}}\cdot\frac{\alpha}{n}&\text{if }a_{i}=1\text{ and }k=0\\ 0&\text{if }a_{i}=1\text{ and }k>0\end{cases},$
where $\beta_{\Omega}$ is the total market power of the lowest-bidding firms. Note that in the heterogeneous case, the set of lowest bidders $\Omega$ is likely to be a singleton, depending on how power parameters are distributed.

### v2

@@ 3.1 The Minimum Price Game: A Single-Stage Formulation
All players are symmetric in their strategy set. Indeed, each agent $i$ chooses a strategy $s_{i}\in S$ where $S=\{\textit{FP},\textit{CP}\}$. The strategy Fair Price (FP) is associated with bidding the fair price $b_{i}$, and Collusive Price (CP) with bidding a higher price, defined by $\alpha\cdot b_{i}$, where $\tau\geq\alpha>1$, with the upper bound $\tau$ ensuring realistic bids. We denote the strategy profile for a given occurrence as the tuple $(s_{1},s_{2},\dots,s_{n})$.
Heterogeneity.
Each agent $i$ has a market power defined by the parameter $\beta_{i}\in(0,1)$, with $\sum_{i\in\mathcal{N}}\beta_{i}=1$, indicating that the average market strength, $\mu(\beta)$, is always $\frac{1}{n}$. This market power might represent the size of the agent or its market share, influencing its ability to reduce costs and leverage economies of scale. All agents share the same cost function, which determines their bids
$b_{i}=(1-\beta_{i})v\quad\forall i\in\mathcal{N},$ (1)
The notion of strong and weak agents is relative to the distribution of $\beta$, but according to (1), the closer $\beta$ is to 1, the lower the agent can bid, thus the stronger it is. Market heterogeneity is quantified by $\sigma(\beta)$, the standard deviation of the distribution of power parameter values. In a perfectly homogeneous market, $\sigma(\beta)=0$ and $\beta_{i}=\frac{1}{n}$ for all $i$.
Remark 1.
As discussed earlier, we need to model heterogeneous agents within a game formulation that incorporates a social dilemma. Additionally, the payoff structure must reflect a reward function that enables adaptive agents to learn their strategies in repeated settings. Example 1 illustrates why a strict application of the minimum price rule is unsuitable in this context. The core issue is that a strong player could dominate any auction regardless of its strategy, undermining the coexistence of heterogeneous agents.
Example 1 (Hard minimum price rule).
Let $b_{weak}$ and $b_{strong}$ be the bids in an auction within a heterogeneous duopoly. Naturally, and according to (1), we have $b_{strong}<b_{weak}$. Assuming the payoffs are determined by a minimum price rule, i.e., $u_{i}=b_{i}$ if $b_{i}=\min(b_{weak},b_{strong})$ and $u_{i}=0$ otherwise, we obtain the following bimatrix game:
[…]
We observe that the outcomes of the asymmetric strategy profiles depend on the level of heterogeneity. The strong agent could potentially win regardless of the strategy profile, while the weak agent might struggle to learn a long-term strategy since it may only receive a payoff of 00 as a reward signal in a repeated setting.
Payoffs.
Let $u_{i}(\textit{FP},k)$ represent the payoff of agent $i$ when it bids the fair price and $k$ opponents also bid the fair price, and let $u_{i}(\textit{CP},k)$ denote the payoff of agent $i$ when it bids the collusive price while $k$ opponents bid the fair price, where $k\in[0,n-1]$. The key idea from the minimum price rule is that coordination must be unanimous to achieve collusive profits. Therefore, the individual payoff for playing CP when all players cooperate surpasses the individual payoff from universal defection, i.e.,
$u_{i}(\textit{CP},k=0)>u_{i}(\textit{FP},k=n-1).$
However, defection must always yield a higher payoff than cooperation when at least one opponent defects, leading to
$u_{i}(\textit{FP},k>0)>u_{i}(\textit{CP},k>0).$
In fact, we can set $u_{i}(\textit{CP},k>0)=0$, meaning that when some players defect (FP) while others cooperate (CP), the cooperating agents receive no payoff, while defecting agents earn a share based on their market strengths relative to the defecting opponents. Let $\Omega\subseteq\mathcal{N}$ denote the set of agents playing FP, and let the total market power of the defecting agents be $\beta_{\Omega}=\sum_{j\in\Omega}\beta_{j}$. Assuming symmetric play where $\beta_{\Omega}=1$, the payoffs in the general heterogeneous case are given in Table 1.
Table 1: MPG Payoffs
Strategy Profile $\mathbf{k=0}$ $\mathbf{k>0}$ $\mathbf{k=n-1}$
$u_{i}(\textit{FP},k)$ $b_{i}$ $\frac{\beta_{i}}{\beta_{\Omega}}\cdot b_{i}$ $\beta_{i}\cdot b_{i}$
$u_{i}(\textit{CP},k)$ $\alpha\cdot\beta_{i}\cdot b_{i}$ 00 00
Analysis.
The parameter $\beta$ controls the incentive to defect. As we can derive from (1) and Table 1, and since $\alpha$ is the collusive factor common to all agents, the incentive factor for agent $i$ associated with defecting while every opponent colludes (FP, $k=0$) is
$\gamma_{i}=\frac{1}{\beta_{i}}.$ (2)
The pairwise relation in terms of incentive factors between agent $i$ and agent $o$ is given by
$\gamma_{i}=\frac{\beta_{o}}{\beta_{i}}\gamma_{o}.$ (3)
The incentive to defect is proportional to the market power $\beta$, and therefore weak agents have more incentive to defect than strong ones, as expected in real-life scenarios [13, 47]. Example 2 proposes a numerical depiction of the incentive to defect in the MPG.
This sharing mechanism allows for heterogeneous agents to coexist while preserving the dynamic stemming from the minimum price. In fact, the power parameter $\beta$ not only represents the ability to cut costs but also the work capacity. As strong agents are more rewarded than weak ones in symmetrical plays (see Table 1), the payoff structure of the MPG emulates the natural turnover in auction participation of heterogeneous agents observed in markets, making the coexistence between weak and strong agents possible. Indeed, a higher payoff mathematically represents the higher frequency of wins by stronger agents in a strict minimum price-ruled process. When the setting simplifies to a homogeneous game, the sharing mechanism in symmetrical plays smooths out a random turnover between agents of equal power, as shown in Example 3.
Example 2 (Incentive factor).
Consider three heterogeneous agents $\{1,2,3\}$, and their associated power parameters $[0.25,0.25,0.5]$. For simplicity, we set $v$ to $1$ in equation (1). In the collusive scenario, all agents choose to play CP, meaning that $\beta_{\Omega}=1$, and according to (1) and Table 1, we have $u_{1}=u_{2}=0.19\alpha$ and $u_{3}=0.25\alpha$, with the strongest agent receiving the highest payoff.
Now consider that agent $1$, a weak agent, chooses to defect instead. In this case, we have $\beta_{\Omega}=0.25$, and thus, $u_{1}=0.75$ and $u_{2}=u_{3}=0$. As predicted by (2), the incentive factor for the weak agent to defect is $4$.
If the strong agent decides to defect while others cooperate, then $\beta_{\Omega}=0.5$, and thus, $u_{1}=u_{2}=0$ and $u_{3}=0.5$. Here, the incentive factor for the strong agent to defect is $2$. As predicted by (3), the incentive factor for the weak agent to defect is twice as high because its market power ($\beta_{1}=\beta_{2}=0.25$) is half that of its stronger opponent.
Example 3 (Homogeneous special case).
Consider a homogeneous duopoly ($\sigma(\beta)=0$ and $n=2$). According to (1) and Table 1, we get the following bimatrix game:
[…]
In symmetrical plays, the payoffs are the same for both agents. This is mathematically equivalent to a random selection mechanism for breaking ties between the lowest bidders under the hard minimum price rule. Since homogeneous players produce the same bid, $b$, a selection mechanism would involve a random process to break ties in symmetrical play. This could be done by either using the uniform distribution, $i^{*}\sim\mathcal{U}(\mathcal{N})$, or by adding a stochastic perturbation to the bids, $b_{i}+\epsilon_{i}$.
@@
The $n$-player homogeneous MPG is a Prisoner’s Dilemma.
@@
Consider a reference player $i\in\mathcal{N}$. Also, consider $\alpha\leq 2$, a realistic upper bound for the collusive factor. The following payoff structure unfolds by definition:
  • $T$: Temptation payoff (player $i$ bids the fair price while all opponents bid their collusive price),
  • $R$: Reward payoff (everybody cooperates),
  • $P$: Punishment payoff (everybody defects),
  • $S$: Sucker’s payoff (player $i$ bids the collusive price while at least one opponent defects).
  1. i.
    $P$ is the dominant strategy profile.
    In the homogeneous MPG, $\beta_{i}=\frac{1}{n}$ for all $i\in\mathcal{N}$, meaning that $b_{i}=b$ for all $i\in\mathcal{N}$ according to (1). Hence,
    $P=u_{i}(\textit{FP},k=n-1)=\frac{b}{n}\quad\text{and}\quad S=u_{i}(\textit{CP},k>0)=0,$
    and $P>S$. Since
    $u_{i}(\textit{FP},k=0)=b>\frac{\alpha b}{n}=u_{i}(\textit{CP},k=0)\quad\forall i\in\mathcal{N}$
    for any $\alpha<2$ (since $n\geq 2$), $T>S$, meaning that FP (defection) is the best response for every player. The strategy profile $P$ is therefore the unique Nash equilibrium.
  2. ii.
    The strategy profile $R$ is Pareto Optimal.
    A strategy profile is Pareto Optimal if no player can improve their outcome without making at least one other player worse off. Since
    $u_{i}(\textit{CP},k=0)=\frac{\alpha b}{n}>\frac{b}{n}=u_{j}(\textit{FP},k=n-1)\quad\forall(i,j)\in\mathcal{N},$
    because $\alpha>1$, $R>P$, and $(\textit{CP},k=0)$ is therefore Pareto Optimal.
  3. iii.
    The payoff structure satisfies the inequality $T>R>P>S$.
    Since
    $u_{i}(\textit{FP},k=0)=b>\frac{\alpha b}{n}=u_{i}(\textit{CP},k=0)\quad\forall i\in\mathcal{N},$
    because $\frac{\alpha}{n}<1$, $T>R$. By combining the results of i. and ii., we conclude $T>R>P>S$.
@@
The $n$-player heterogeneous MPG is not a Prisoner’s Dilemma.
@@
In the heterogeneous MPG, the market powers differ among agents, leading to variations in the payoffs. If the level of heterogeneity is high enough, the strongest agent may receive a higher payoff with $R$ (collusion) than with $T$ (being the only defector). In this case,
$u_{i}(\textit{CP},k=0)=\alpha\beta_{i}b_{i}\quad\forall i\in\mathcal{N},$
if $\alpha\beta_{i}>1$, then $u_{i}(\textit{CP},k=0)=R>T=u_{i}(\textit{FP},k=0)$, and condition iii. in the proof of Proposition 1 no longer holds. ∎
@@ 3.2 Markov Game Formulation
Definition 1.
Markov game. A Markov game, also known as a stochastic game, is a generalization of a Markov Decision Process (MDP) to multiple interacting agents contained in the set $N$. Each player $i$ has an associated action space $\mathcal{A}_{i}$. The game is defined over a common state space $\mathcal{S}$. The dynamics of the game are captured by a transition function $T:\left(\bigtimes_{i\in N}\mathcal{A}_{i}\right)\times\mathcal{S}\rightarrow[0,1]$, which determines the probability of moving from one state to another based on the actions of all players. Each player $i$ has a reward function $R_{i}:\left(\bigtimes_{i\in N}\mathcal{A}_{i}\right)\times\mathcal{S}\rightarrow\mathbb{R}$, defining the payoff $r_{i}$ for player $i$ based on the current state and actions taken by all players. The objective for each player $i$ is to maximize the expected sum of discounted rewards, expressed as $E\left[\sum_{k=0}^{\infty}\gamma^{k}r_{i,t+k}\right]$, where $\gamma$ is the discount factor that values present rewards over future rewards. A key feature of an MDP, and therefore of a Markov game, is the Markov property, which asserts that future states depend only on the current state and the actions taken by the agents, not on the sequence of events that preceded it, ensuring that past states are irrelevant once the current state is known.
Action space.
The action space $\mathcal{A}=\{0,1\}$ is binary, reflecting the two strategies available to players, with bidding the fair price being represented by 00 and bidding the collusive price with $1$. This representation simplifies the computation of action and joint action frequencies. Let $a_{i}^{t}\in\mathcal{A}$ be the action taken by agent $i$ at time $t$, and $f^{t}\in\mathcal{A}^{n}$ be the joint action $(a_{1},\dots,a_{n})$ in auction $t$. The action and joint action frequencies up to iteration $t$ are then respectively given by
@@
State space.
The choice of state space in a Markov game is often at the discretion of the implementer, depending on the specifics of the problem and the goals of the implementation. However, for our purposes, we define the MPMG with core variables that capture essential aspects of the game. Specifically, these sets include all players’ information on action frequencies $\mathbf{\bar{a}}^{t}=\{\bar{a}_{j}^{t}\}_{j\in\mathcal{N}}$, joint action frequencies $\mathbf{\bar{f}}^{t}=\{\bar{f}_{j}^{t}\}_{j\in\mathcal{N}}$, average rewards $\mathbf{\bar{r}}^{t}=\{\bar{r}_{j}^{t}\}_{j\in\mathcal{N}}$, and power parameters $\mathbf{\beta}=\{\beta_{j}\}_{j\in\mathcal{N}}$. Therefore, the state space at any time $t$ is $\mathcal{S}_{t}=\{\mathbf{\bar{a}}^{t},\mathbf{\bar{f}}^{t},\mathbf{\bar{r}}^{t},\mathbf{\beta}\}$, and is of size $2n+|\mathcal{A}|^{n}$. This minimal configuration for the state space aligns with Assumption 4. Furthermore, the power parameters also allow for the possibility of modeling a dynamic market in which agents could lose or gain market power.
Reward function.
The payoff structure defined for the MPG translates to a reward function in the MPMG, emphasizing the inherent connection between the Markov game and its RL implementation. The reward function, together with the transition function, which defines how the game state evolves in response to actions, fundamentally shapes the dynamics and outcomes of the game. In fact, in the classical RL setting, the normalized reward function for agent $i$ at each iteration $t$ is expressed as
@@
where $\gamma$ is the discount factor and $r_{i}^{t-m}$ is the immediate reward received by agent $i$ at time $t-m$. The MPMG relies on the tension between Nash and Pareto dominant strategies inherent in the classical Prisoner’s Dilemma. However, as with the state space, the reward function can be adjusted for specific implementation purposes (e.g., reward shaping) as long as the fundamental dynamics are preserved. Following Assumption 3, we can further assume that the contract value $v$ remains equivalent as the iterations unfold, normalizing it to $1$. Therefore, (1) becomes
@@
We do not make this assumption formal, as this normalization does not affect the generality of the model but only simplifies the calculations. The immediate rewards for agent $i$ at nay time $t$ are given in table 1.
Assumption 5 (Single Public Contractor).
All public contractors share the same goal.
In real-world scenarios, the set of public contractors in a given procurement market usually includes many entities. Despite this plurality, it is safe to assume that all public contractors work towards the same overarching goal. Assumption 5 allows us to reduce the set of public contractors to a singleton.
Dynamic.
The MPMG unfolds as an episodic game of arbitrary length (number of episodes), with each episode consisting of a single-step auction (Assumptions 1 and 2). Agents compete against each other in an environment representing the same entity in each episode (Assumption 5). Due to the Markov property, memory is encapsulated in the observation space $\mathcal{S}$ at any episode $t$. Agents use this observation space to learn their strategies, conditional on occurrences in $\mathcal{S}$, effectively playing a Markov game where the reward function, $R(a_{1},\dots,a_{n})$, is deterministic and based on the joint action. The transition probability function is given by $P(s^{\prime}|s,a)$, where $s^{\prime}$ is the new state and $s\in\mathcal{S}$. When the state space is not used for decision-making, the agents are effectively playing an iterated MPG.